OpenStudy (anonymous):

Evaluate (ab) 2 for a = 4 and b= 3. A. 81 B. 144 C. 24 D. 36 i think its b

Vocaloid (vocaloid):

is the 2 supposed to be an exponent?

OpenStudy (anonymous):

Okay so if we have (ab)2 and a=4 and b=3 then we need to replace a and b with there values. Now before we do that we are going to set up our equation. Whenever you have things like "Ab" or "5x" you multiply them. So "A*B In this case (a*b)2 and since it doesn't say weather to add or subtract or anything like that, we multiply *2 So, (4*3)*2 4*3=12 12*2=24 Final answer: 24.

OpenStudy (anonymous):

^That is if the 2 were not to be a exponent and rather just a value.

OpenStudy (anonymous):

And if the 2 is a exponent we do the same thing, but this time you solve as you would with a exponent. So we have: \[(4*3)^{2}\] And so we do 4*3 4*3=12 Now we have \[12^{2}\] And to solve we just do 12*12 12*12=144 If you didn't know if it was a exponent, if it is it will be a small to in the top right of the equation.

OpenStudy (anonymous):

@Vocaloid is this correct?